OpenStudy (anonymous):

eight years ago. a man was four years less than three times as old as his son's age.six years from now.he will be twice as old as his son.how old are they now?

jigglypuff314 (jigglypuff314):

(M-8) = 3(S-8) - 4 (M+6) = 2(S+6)

jigglypuff314 (jigglypuff314):

M stands for the Man's age presently S stands for the man's Son's age presently "eight years ago" so minusing 8 from both the man and the son in the first equation "six years from now" so adding 6 to both the man and the son in the second equation

OpenStudy (anonymous):

can you give me the complete solutions?

jigglypuff314 (jigglypuff314):

how about we try to solve it together? simplify the equations using distributive property (M-8) = 3(S-8) - 4 (M+6) = 2(S+6)

OpenStudy (anonymous):

thanks a lot :D

jigglypuff314 (jigglypuff314):

so (M-8) = 3(S-8) - 4 becomes M-8 = 3S - 24 - 4 and (M+6) = 2(S+6) becomes M+6 = 2S + 12

OpenStudy (anonymous):

then m=2s+6?

jigglypuff314 (jigglypuff314):

yes! so then since M = M you can plug in what you got into the first equation so 2s+6 = 3s+20 solve for s

OpenStudy (anonymous):

the answer will be negative?

jigglypuff314 (jigglypuff314):

oh sorry! typo 2s+6 = 3s-20

OpenStudy (anonymous):

m-8? where is the 8?

jigglypuff314 (jigglypuff314):

carried it over 2s+6 - 8 = 3s - 24 - 4 2s + 6 = 3s - 28 + 8 2s + 6 = 3s - 20

OpenStudy (anonymous):

how can i compute m?

jigglypuff314 (jigglypuff314):

first find s then you can plug s back into one of the original equations to get m :)

OpenStudy (anonymous):

thanks a lot

jigglypuff314 (jigglypuff314):

glad I can help :)